# Startup Acquisitions: Acquihires and Talent Hoarding

### v3

@@ B.1 People and Technology
Suppose firm 1 has not made the acquisition. It follows from Assumption B1 that firm 2 makes the acquisition if and only if it has a high match type, irrespective of its beliefs. Moving to stage 1, firm 1’s beliefs are given by the prior belief. Suppose a high-match firm 1 acquired the startup. By Assumption B2 a low-match firm 2 would not buy the technology part of the startup and by our tie-breaking assumption neither would a high-match firm 2. Then, by Assumption B1 a high-match firm acquires the startup and keeps the technology. Suppose a low-match firm 1 acquired the startup. By our tie-breaking assumption, a low-match firm 2 would not buy the startup’s technology. However, by Assumption B2, the technology part would be sold to a high-match firm 2. Anticipating this, the threshold for a low-match firm 1 to acquire the startup changes relative to the model in the main text. Formally, doing nothing yields
@@
$\Pi_{F}-\pi_{E}+(1-\lambda)\bar{\pi}_{F}^{L}+\lambda(q+(1-\delta)\bar{\pi}_{F}^{L}-\delta\underline{\pi}_{F}^{H}),$
@@
$q=\frac{\delta\left(\bar{\pi}_{F}^{H}+\underline{\pi}_{F}^{L}+\bar{\pi}_{F}^{L}+\underline{\pi}_{F}^{H}\right)}{2},$
@@
$\Pi_{F}-\pi_{E}+(1-\lambda)\bar{\pi}_{F}^{L}+\lambda\left(\frac{\delta\left(\bar{\pi}_{F}^{H}+\underline{\pi}_{F}^{L}+\bar{\pi}_{F}^{L}+\underline{\pi}_{F}^{H}\right)}{2}+(1-\delta)\bar{\pi}_{F}^{L}-\delta\underline{\pi}_{F}^{H}\right)\geq\Pi_{F}-\lambda\underline{\pi}_{F}^{H},$
@@
One can verify that for $\delta=0$ the above condition reduces to the condition (1) in the main text. As $\delta$ increases, the threshold $\lambda_{A}(\delta)$ decreases, i.e., the acquisition happens for a larger set of parameters. Intuitively, as the technology part of the startup can be sold to a high-match firm 2, the expected cost of hoarding the talent falls for the low-match firm 1, leading to more talent hoarding.
B.2 Dominant Firm
@@ B.3 Multiple Firms
B.4 Partial Acquisitions
This section extends the baseline model to incorporate the possibility of partial investments. That is, a firm may acquire a (minority) stake in the startup without integrating it. To do so, we need to specify the payoffs resulting from such partial acquisitions as well as the rights that come with partial ownership.
Formally, the ability to make partial acquisitions means that firms can additionally try to acquire a share $s\in(0,1]$ of the startup and continue to operate it as a stand-alone entity. In contrast to the model in the main text, we allow for upfront and deferred payments $(p,d)$ as is standard in such transactions. In the main text, doing so would not change anything. In the presence of partial ownership, it allows the acquiring firm to distinguish at least partially between the investor (who only gets a part of the upfront payment) and the entrepreneur (who can also receive deferred payments). If a firm acquires a share $s$ with bid $(p,d)$, its payoff reads $\Pi_{F}+s\pi_{E}(s)-p-d$, while the entrepreneur’s payoff is $(1-s)\pi_{E}(s)+w(s)+p+d$. Here, $\pi_{E}(s)$ captures the startup’s profit net of potential wages paid to the entrepreneur as a function of the size of the external ownership. These profits accrue to the firm and the entrepreneur proportionate to their stake in the startup. Correspondingly, $w(s)$ constitutes the entrepreneur’s wage (net of effort costs) for different degrees of outside ownership. In particular, $\pi_{E}=\pi_{E}(0)+w(0)$. That is, when the entrepreneur owns the entire startup and thus obtains all of its profit and “pays herself” a net-of-effort wage, her payoff coincides with the initial payoff in the baseline model. The other firm’s payoff is unchanged at $\Pi_{F}$. When a firm acquires a stake in the startup it receives a share of profits, while the dilution of ownership gives rise to moral hazard on the entrepreneur’s side, which we capture in reduced form, only imposing the following assumption.
Assumption B3.
We assume that $\pi_{E}(s)+w(s)$ is decreasing in $s$ with $\pi_{E}(0)+w(0)>\pi_{E}(1)+b(1)$.
Assumption B3 captures the moral hazard arising when the entrepreneur no longer fully owns the startup. $\pi_{E}(s)+w(s)$ being decreasing reflects the reduced effort of the entrepreneur as a result of agency. $\pi_{E}(0)+w(0)>\pi_{E}(1)+w(1)$ captures that the first-best value of a startup fully owned by the entrepreneur is strictly higher than the second-best value, when the entrepreneur is actually an employee and has no stake in the startup.
The timing of the game is as follows. In the first stage, firm 1 has the opportunity to make a bid to the entrepreneur to acquire a share $s_{1}\in(0,1]$ of the startup or make an acquihire. If the entrepreneur accepts a bid for an acquihire, the game ends. Otherwise, we move to stage 2, the ownership structure of the startup depending on whether the entrepreneur accepted firm 1’s bid. In the second stage, firm 2 can make an acquihire by making a (per-share) bid to the owner(s) of the startup, where we restrict attention to bids $(p,d)$ with $p\geq\pi_{E}(s_{1})$, so that no owner can be expropriated. If the entrepreneur does not accept the bid, the game ends. If the entrepreneur accepts, firm 1 can either also accept or try to block the transaction and succeeds in doing so with probability $q(s_{1})$, where $q$ is a weakly increasing function with $q(0)=0$ and $q(1)=1$. Nature determines whether a potential blocking attempt succeeds and the game ends.
Proposition B3.
Under Assumptions 1 and B3 and without blocking rights, i.e., $q(s)=0$ for all $s\in[0,1]$, firm 1’s behavior in any PBE is uniquely specified. Namely, if firm 1 draws a high match type, it will make an acquihire; if it draws a low type it will make an acquihire if and only if $\lambda\geq\lambda_{A}$ and do nothing otherwise.
Proof.
See the proof of Proposition B4, which contains this as a special case.
The result in Proposition B3 shows that partial ownership of the startup is not enough to change firm 1’s behavior relative to the setting with only acquihires in Proposition 1. Indeed, since an investment is not profitable in itself and does not prevent a high-match firm 2 from making an acquihire, firm 1 will continue to either do nothing or make an acquihire, depending on the probability of a high-match firm 2 materializing. As the next result shows, it is ownership accompanied by some measure of control over the startup, which makes investments attractive to firm 1.
Proposition B4.
Under Assumptions 1 and B3 and with blocking rights, firm 1’s behavior in any PBE is uniquely specified. Namely, if firm 1 draws a high type, it will make an acquihire. If it draws a low type, there is a threshold value for $\lambda$ below which it does nothing. Above the threshold, it will do an investment and, depending on the model’s parameters, there may be an even higher threshold above which it does an acquihire.
Comparing the thresholds across Propositions 1, B3 and B4, we can observe that the presence of blocking rights on investments increases the parameter space for which some form of talent hoarding occurs. However, the possibility of making investments instead of an acquihire may also have “mitigating” effects relative to a setting with only acquihires, as talent hoarding by means of investments create smaller inefficiencies than by means of acquihires. Put differently, allowing for investments increases the extensive margin of talent hoarding but partially decreases its intensive margin.
Proof.
We solve the game backward.
Stage 2
[…]
Hence, firm 1 would try to block any bid $(p_{2},d_{2})$ resulting in a lower payoff than the above, while the entrepreneur would not accept any bid yielding a lower payoff than the above. Given these constraints, firm 2 will choose among three options: (i) do nothing and receive payoff $\Pi_{F}$; (ii) make the minimum bid such that both firm 1 and the entrepreneur accept, yielding payoff $\bar{\Pi}_{F}^{\theta}-\pi_{E}(s_{1})-w(s_{1})-\Pi_{F}+\underline{\Pi}_{F}^{\theta}$ for firm 2; or (iii) make a bid that only the entrepreneur accepts, risking a blocking attempt by firm 1. This third option provides an expected payoff for firm 2 of $q(s_{1})\Pi_{F}+(1-q(s_{1}))(\bar{\Pi}_{F}^{\theta}-\pi_{E}(s_{1})-w(s_{1}))$.
$\bar{\Pi}_{F}^{L}-\pi_{E}(s_{1})-w(s_{1})\geq\Pi_{F}.$
It follows from Assumptions 1 and B3 that this condition is not necessarily satisfied. Let $\hat{s}$ be the threshold above which this condition is satisfied. Now let’s compare doing nothing with inducing both for $s_{1}<\hat{s}$. We obtain that inducing both is better whenever
$\bar{\Pi}_{F}^{L}-\pi_{E}(s_{1})-w(s_{1})-\Pi_{F}+\underline{\Pi}_{F}^{L}\geq q(s_{1})\Pi_{F}+(1-q(s_{1}))(\bar{\Pi}_{F}^{L}-\pi_{E}(s_{1})-w(s_{1})),$
which, after rearrangement, can be rewritten as
$q(s_{1})(\bar{\Pi}_{F}^{L}-\Pi_{F}-\pi_{E}(s_{1})-w(s_{1}))\geq\Pi_{F}-\underline{\Pi}_{F}^{L}$ (B.8)
Since the right-hand side of (B.8) is positive and the left-hand side of it is negative for $s_{1}<\hat{s}$, this condition is never satisfied. Thus, for $s_{1}<\hat{s}$ the low type does nothing.
Now let us compare firm 2 inducing only the entrepreneur versus inducing both to accept when $s_{1}\geq\hat{s}$. As above, inducing both is better whenever condition (B.8) holds. Since the left-hand side of (B.8) is increasing in $s_{1}$, we define $s^{L}\in[\hat{s},1]$ as the threshold above which the condition is satisfied. Overall, a low-match firm 2 will take the following actions: for $s_{1}<\hat{s}$, do nothing; for $\hat{s}\leq s_{1}\leq s_{L}$, induces only the entrepreneur to accept; and for $s_{1}>s_{L}$, induce both the entrepreneur and firm $1$ to accept the offer.
$q(s_{1})(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E}(s_{1})-w(s_{1}))\geq\Pi_{F}-\underline{\Pi}_{F}^{H},$
Stage 1
$(N,E):$ $\;\lambda(1-(q(s_{1})))\underline{\Pi}_{F}^{H}+(1-\lambda(1-q(s_{1})))\Pi_{F}+\Delta(s_{1})$
$(N,B):$ $\;\Pi_{F}+\Delta(s_{1})$
$(E,B):$ $\;(1-(1-\lambda)(1-q(s_{1})))\Pi_{F}+(1-\lambda)(1-q(s_{1}))\underline{\Pi}_{F}^{L}+\Delta(s_{1})$
$(B,B):$ $\;\Pi_{F}+\Delta(s_{1})$
$(E,E):$ $\;q(s_{1})\Pi_{F}+(1-q(s_{1}))(\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\underline{\Pi}_{F}^{L})+\Delta(s_{1})$
$(B,E):$ $\;\lambda(1-(q(s_{1})))\underline{\Pi}_{F}^{H}+(1-\lambda(1-q(s_{1})))\Pi_{F}+\Delta(s_{1})$
$\Pi_{F}+s_{1}\pi_{E}(s_{1})-(\pi_{E}(0)+w(0)-(1-s_{1})\pi_{E}(s_{1})-w(s_{1}))=\Pi_{F}+\Delta(s_{1}).$ (B.9)
A high-match firm 2 would make a bid that firm 1 will try to block, succeeding with probability $q(s_{1})$, yielding the following payoff to firm $1$
$q(s_{1})\left(\Pi_{F}+\Delta(s_{1})\right)+(1-q(s_{1}))\left(\underline{\Pi}_{F}^{H}+\Delta(s_{1})\right).$ (B.10)
Adding (B.9) and (B.10) up while multiplying them with the probabilities $1-\lambda$ and $\lambda$, respectively, we obtain the expression in the above list. [^16] Observe that the size of firm 1’s investment $s_{1}$ is not the same across cases in the above list, as different investment sizes induce the different behaviors of firm 2. Finally, to complete the list, note that an acquihire gives a payoff $\bar{\Pi}_{F}^{L}-\pi_{E}(0)-w(0)$ to firm $1$ and doing nothing results in $\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\Pi_{F}$.
To determine firm 1’s equilibrium strategy, we need to further distinguish between three cases three cases based on the values of the thresholds $s^{H}$, $s^{L}$, and $\hat{s}$. Recall that a low-match firm 2 will do nothing below $\hat{s}$, induce the entrepreneur between $\hat{s}$ and $s^{L}$ and potentially induce both above $s^{L}$, while a high-match firm 2 will induce only the entrepreneur below $s^{H}$ and may induce both above it.
$(N,E):$ $\;\lambda(1-(q(s_{1})))\underline{\Pi}_{F}^{H}+(1-\lambda(1-q(s_{1})))\Pi_{F}+\Delta(s_{1})$
$(N,B):$ $\;\Pi_{F}+\Delta(s_{1})$
$(N):$ $\;\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\Pi_{F},$
B.5 Unknown Order of Moves
@@ B.6 Surplus Sharing
We solve the game backward. Consider a high-match firm 2. The surplus resulting from an acquihire is given by $\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E}>0$ so that an acquihire takes place and the resulting payoffs for firm 2 and the entrepreneur read $\Pi_{F}+(1-\gamma)(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})$ and $\pi_{E}+\gamma(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})$, respectively. Consider a low-match firm 2. The surplus then reads $\bar{\Pi}_{F}^{L}-\Pi_{F}-\pi_{E}<0$ so that no acquihire takes place. Moving to period 1, consider a high-match firm 1. The surplus resulting from an acquihire reads

### v4

@@ A.2 Proof of Proposition 4
Suppose firm 1 has not made the acquisition. It follows from Assumption 3 that firm 2 makes the acquisition if and only if it has a high match type, irrespective of its beliefs. Moving to stage 1, firm 1’s beliefs are given by the prior belief. Suppose a high-match firm 1 acquired the startup. By Assumption 4 a low-match firm 2 would not buy the technology part of the startup and by our tie-breaking assumption neither would a high-match firm 2. Then, by Assumption 3 a high-match firm acquires the startup and keeps the technology. Suppose a low-match firm 1 acquired the startup. By our tie-breaking assumption, a low-match firm 2 would not buy the startup’s technology. However, by Assumption 4, the technology part would be sold to a high-match firm 2. Anticipating this, the threshold for a low-match firm 1 to acquire the startup changes relative to the model in the main text. Formally, doing nothing yields
@@
$\Pi_{F}-\pi_{E}+(1-\lambda)\bar{\pi}_{F}^{L}+\lambda(q+(1-\tau)\bar{\pi}_{F}^{L}-\tau\underline{\pi}_{F}^{H}),$
@@
$q=\frac{\tau\left(\bar{\pi}_{F}^{H}+\underline{\pi}_{F}^{L}+\bar{\pi}_{F}^{L}+\underline{\pi}_{F}^{H}\right)}{2},$
@@
$\Pi_{F}-\pi_{E}+(1-\lambda)\bar{\pi}_{F}^{L}+\lambda\left(\frac{\tau\left(\bar{\pi}_{F}^{H}+\underline{\pi}_{F}^{L}+\bar{\pi}_{F}^{L}+\underline{\pi}_{F}^{H}\right)}{2}+(1-\tau)\bar{\pi}_{F}^{L}-\tau\underline{\pi}_{F}^{H}\right)\geq\Pi_{F}-\lambda\underline{\pi}_{F}^{H},$
@@
A.3 Proofs of Propositions 5 and 6
In what follows, we prove Proposition 6, which is a generalization of Proposition 5. We solve the game backward.
Stage 2
[…]
Hence, firm 1 would try to block any bid $(p_{2},d_{2})$ resulting in a lower payoff than the above, while the entrepreneur would not accept any bid yielding a lower payoff than the above. Given these constraints, firm 2 will choose among three options: (i) do nothing and receive payoff $\Pi_{F}$; (ii) make the minimum bid such that both firm 1 and the entrepreneur accept, yielding payoff $\bar{\Pi}_{F}^{\theta}-\pi_{E}(s_{1})-w(s_{1})-\Pi_{F}+\underline{\Pi}_{F}^{\theta}$ for firm 2; or (iii) make a bid that only the entrepreneur accepts, risking a blocking attempt by firm 1. This third option provides an expected payoff for firm 2 of $\beta(s_{1})\Pi_{F}+(1-\beta(s_{1}))(\bar{\Pi}_{F}^{\theta}-\pi_{E}(s_{1})-w(s_{1}))$.
$\bar{\Pi}_{F}^{L}-\pi_{E}(s_{1})-w(s_{1})\geq\Pi_{F}.$
It follows from Assumptions 1 and 5 that this condition is not necessarily satisfied. Let $\hat{s}$ be the threshold above which this condition is satisfied. Now let’s compare doing nothing with inducing both for $s_{1}<\hat{s}$. We obtain that inducing both is better whenever
$\bar{\Pi}_{F}^{L}-\pi_{E}(s_{1})-w(s_{1})-\Pi_{F}+\underline{\Pi}_{F}^{L}\geq\beta(s_{1})\Pi_{F}+(1-\beta(s_{1}))(\bar{\Pi}_{F}^{L}-\pi_{E}(s_{1})-w(s_{1})),$
which, after rearrangement, can be rewritten as
$\beta(s_{1})(\bar{\Pi}_{F}^{L}-\Pi_{F}-\pi_{E}(s_{1})-w(s_{1}))\geq\Pi_{F}-\underline{\Pi}_{F}^{L}.$ (A.3)
Since the right-hand side of (A.3) is positive and the left-hand side of it is negative for $s_{1}<\hat{s}$, this condition is never satisfied. Thus, for $s_{1}<\hat{s}$ the low type does nothing.
Now let us compare firm 2 inducing only the entrepreneur versus inducing both to accept when $s_{1}\geq\hat{s}$. As above, inducing both is better whenever condition (A.3) holds. Since the left-hand side of (A.3) is increasing in $s_{1}$, we define $s^{L}\in[\hat{s},1]$ as the threshold above which the condition is satisfied. Overall, a low-match firm 2 will take the following actions: for $s_{1}<\hat{s}$, do nothing; for $\hat{s}\leq s_{1}\leq s_{L}$, induces only the entrepreneur to accept; and for $s_{1}>s_{L}$, induce both the entrepreneur and firm $1$ to accept the offer.
$\beta(s_{1})(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E}(s_{1})-w(s_{1}))\geq\Pi_{F}-\underline{\Pi}_{F}^{H},$
Stage 1
$(N,E):$ $\;\lambda(1-(\beta(s_{1})))\underline{\Pi}_{F}^{H}+(1-\lambda(1-\beta(s_{1})))\Pi_{F}+\Delta(s_{1})$
$(N,B):$ $\;\Pi_{F}+\Delta(s_{1})$
$(E,B):$ $\;(1-(1-\lambda)(1-\beta(s_{1})))\Pi_{F}+(1-\lambda)(1-\beta(s_{1}))\underline{\Pi}_{F}^{L}+\Delta(s_{1})$
$(B,B):$ $\;\Pi_{F}+\Delta(s_{1})$
$(E,E):$ $\;\beta(s_{1})\Pi_{F}+(1-\beta(s_{1}))(\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\underline{\Pi}_{F}^{L})+\Delta(s_{1})$
$(B,E):$ $\;\lambda(1-(\beta(s_{1})))\underline{\Pi}_{F}^{H}+(1-\lambda(1-\beta(s_{1})))\Pi_{F}+\Delta(s_{1})$
$\Pi_{F}+s_{1}\pi_{E}(s_{1})-(\pi_{E}(0)+w(0)-(1-s_{1})\pi_{E}(s_{1})-w(s_{1}))=\Pi_{F}+\Delta(s_{1}).$ (A.4)
A high-match firm 2 would make a bid that firm 1 will try to block, succeeding with probability $\beta(s_{1})$, yielding the following payoff to firm $1$
$\beta(s_{1})\left(\Pi_{F}+\Delta(s_{1})\right)+(1-\beta(s_{1}))\left(\underline{\Pi}_{F}^{H}+\Delta(s_{1})\right).$ (A.5)
Adding (A.4) and (A.5) up while multiplying them with the probabilities $1-\lambda$ and $\lambda$, respectively, we obtain the expression in the above list. [^18] Observe that the size of firm 1’s investment $s_{1}$ is not the same across cases in the above list, as different investment sizes induce the different behaviors of firm 2. Finally, to complete the list, note that an acquihire gives a payoff $\bar{\Pi}_{F}^{L}-\pi_{E}(0)-w(0)$ to firm $1$ and doing nothing results in $\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\Pi_{F}$.
To determine firm 1’s equilibrium strategy, we need to further distinguish between three cases based on the values of the thresholds $s^{H}$, $s^{L}$, and $\hat{s}$. Recall that a low-match firm 2 will do nothing below $\hat{s}$, induce the entrepreneur between $\hat{s}$ and $s^{L}$ and potentially induce both above $s^{L}$, while a high-match firm 2 will induce only the entrepreneur below $s^{H}$ and may induce both above it.
$(N,E):$ $\;\lambda(1-(\beta(s_{1})))\underline{\Pi}_{F}^{H}+(1-\lambda(1-\beta(s_{1})))\Pi_{F}+\Delta(s_{1})$
$(N,B):$ $\;\Pi_{F}+\Delta(s_{1})$
$(N):$ $\;\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\Pi_{F},$
Appendix B Further Extensions
B.1 Dominant Firm
@@ B.2 Multiple Firms
B.3 Uncertain Order of Moves
@@ B.4 Surplus Sharing
We solve the game backward. Consider a high-match firm 2. The surplus resulting from an acquihire is given by $\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E}>0$ so that an acquihire takes place and the resulting payoffs for firm 2 and the entrepreneur read $\Pi_{F}+(1-\gamma)(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})$ and $\pi_{E}+\gamma(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})$, respectively. Consider a low-match firm 2. The surplus then reads $\bar{\Pi}_{F}^{L}-\Pi_{F}-\pi_{E}<0$ so that no acquihire takes place.
Moving to period 1, consider a high-match firm 1. The surplus resulting from an acquihire reads